OpenStudy (anonymous):

Circles M and K are congruent,arc QRis congruent to arc LN and arc OP is congruent to arc VW. Find x and y.

OpenStudy (anonymous):



OpenStudy (anonymous):

If arc QR is congruent to arc LN, so has the same shape and size, this means that 3x + 4y = 8. If are OP is congruent to arc VW, then 5y + 13 = 2x. You can solve: 3x + 4y = 8 2x - 5y = 13

OpenStudy (anonymous):

so it would be x=7 and y=3

OpenStudy (anonymous):

Does that satisfy the equations?

OpenStudy (anonymous):

The best way to solve these equations is to write one variable in terms of the other for one equation, then filling it in into the other equation :) For example: 3x + 4y = 8 4y = 8 - 3x y = 2 - (3/4)x Try substituting this expression for y into the second equation.

OpenStudy (anonymous):

x=13.14

OpenStudy (anonymous):

You would have the correct answer, but I believe you switched a minus sign somewhere :) 2x - 5*( 2 - (3/4)x) = 2x - 10 + (15/4)x = 13 Multiplying the entire equation by 4 for convenience: 8x - 40 + 15x = 52

OpenStudy (anonymous):

By the way, when you find x, you can substitute it into either equation to find y

OpenStudy (anonymous):

i still get the same answer

OpenStudy (anonymous):

Recheck the brackets and minus signs then please :) Could you show us the substitution step of the first equation into the second?

OpenStudy (anonymous):